OpenStudy (anonymous):

Need Help to solve this differential equation \[\frac{dy}{dx} + \frac{2}{\tan(2x)} \times y = \cos(2x)\]

OpenStudy (anonymous):

its integrating factor method. so the integrating factor here will be \[\sin(2x)\]

OpenStudy (anonymous):

is that right?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

so \[\sin(2x)y = \int\limits{}^{}\frac{\cos(2x)\sin(2x)}{\tan(2x)}\]

OpenStudy (anonymous):

the original equation should have been \[\frac{dy}{dx} + \frac{2}{\tan(2x)} \times y = \frac{\cos(2x)}{\tan(2x)}\]

OpenStudy (turingtest):

where did the tan(2x) come from?

OpenStudy (anonymous):

sorry i typed it wrong.

OpenStudy (anonymous):

the very first equation was \[2y + \frac{dy}{dx}\tan(2x) = \cos(2x)\] so i divided by tan to get it linear

OpenStudy (turingtest):

ok, then so far so good you should simplify \[\frac{\sin(2x)\cos(2x)}{\tan(2x)}\]before integrating though btw my connection is horrible, sorry if I'm slow to reply.

OpenStudy (anonymous):

\[\int\limits_{}^{} \cos^2(2x) dx\]

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

not sure how to do this.

OpenStudy (anonymous):

cos(4x) = 2cos^2(2x) - 1

OpenStudy (anonymous):

so \[1/2 \int\limits_{}^{} 2\cos^2(2x) - 1 + 1 = 1/2 \int\limits_{}^{} \cos(4x) + 1\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

btw I know @Coolsector 's formula better as\[\cos^2x=\frac12(1+\cos(2x))\]just so you know where it came from...

OpenStudy (anonymous):

\[y = \frac{\frac{1}{8} \sin(4x) + x}{\sin(2x)}\] ?

OpenStudy (anonymous):

so using yours Turing would give: \[\int\limits_{}^{} \frac{1}{2} + \frac{1}{2}\cos(2x) = \frac{1}{2}x + \frac{1}{4}\sin(2x)\]

OpenStudy (anonymous):

sorry i was doing cos^2(x) not cos^2(2x)

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} \frac{1}{2} + \frac{1}{2}\cos(2x) = \frac{1}{2}x + \frac{1}{8}\sin(2x)\]

OpenStudy (anonymous):

is that right?

OpenStudy (turingtest):

no, the argument after using the trig formula is 4x \[\cos^2(x)=\frac12(1+\cos(2x))\implies\cos^2(2x)=\frac12(1-\cos(4x))\]so you and coolsector had it right I was just trying to make a point, sorry if I confused you

OpenStudy (anonymous):

yeah sorry i didnt get that.

OpenStudy (turingtest):

typo, that should be a + in the second equation above...

OpenStudy (anonymous):

\[y = \frac{ \sin(4x) }{8\sin(2x)} + \frac{x}{8}\] ?

OpenStudy (anonymous):

ahh i see Turing. So is that the solution for y? (plus a constant)

OpenStudy (turingtest):

\[\int\frac12+\frac12(\cos(4x))dx=\frac12x+\frac18\sin(4x)+C=y\sin(2x)\]differential equations is when you find out how incredibly important the integration constant C is; it will *completely* ruin your answer if you neglect it. solve for y

OpenStudy (anonymous):

always forget C :(

OpenStudy (anonymous):

the problem is an initial value problem y(pi/4) = 1

OpenStudy (turingtest):

ok, first solve for y to get the general you don't have to do it that way first, but it's a good habit

OpenStudy (anonymous):

so sin(4pi/4) = 0 and sin(2pi/4) = 1 which gives \[\frac{\pi}{8} = y\]

OpenStudy (anonymous):

no sorry , pi/8 = -C

OpenStudy (turingtest):

that's right :) your explicit solution really should utilize more trig identities as well

OpenStudy (anonymous):

it does seem quite complicated. Which identities do you think?

OpenStudy (turingtest):

\[\frac12x+\frac18\sin(4x)+C=y\sin(2x)\]using the double angle formula for sine\[\sin(2x)=2\sin x\cos x\]we get\[\frac12x+\frac14\sin(2x)\cos(2x)+C=y\sin (2x)\]then just divide by sin(2x) and remember that 1/sinx=cscx to make it look nicer

OpenStudy (anonymous):

Awesome! Im useless at remembering identities too

OpenStudy (anonymous):

thanks.

OpenStudy (turingtest):

no problem :)